OpenStudy (anonymous):

A small block of mass m1 is placed on a larger block of mass m2. The blocks are placed on an inclined plane with an angle theta. Suppose a known force of magnitude P is applied on that string so that m2 accelerates and m1 slides on m2. If the coefficient of kinetic friction between m1 and m2 is u(k1) and the coefficient of kinetic friction between the m2 and the table is u(kT) what will the acceleration of m2 be?

OpenStudy (anonymous):



OpenStudy (anonymous):

Ok so I am just unsure as to how block m1 effects the acceleration of block m2. I know that (m2)*a=P-(m2)*g*sin(theta)-(m2)*g*u(kT)*cos(theta)

OpenStudy (anonymous):

I think I am supposed to subtract the frictional force of m1 because it acts between m1 and m2 but am not 100% sure. Also is there anything else to add?

OpenStudy (anonymous):

The equation that you derived isn't right

OpenStudy (anonymous):

try drawing a free body diagram for each block and then labelling all forces before writing down the equation

OpenStudy (anonymous):

well that isn't all of the equation

OpenStudy (anonymous):

that is just for the bottom block. I was wondering how forces on m1 will effect block 2

OpenStudy (anonymous):

even if its for the bottom block, its still wrong.

OpenStudy (anonymous):

think about it

OpenStudy (anonymous):

if there wasn't a block on top of m2 then that would be the equation

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so my question is what other forces do I need to add to the equation because of m1 sliding on top of m2

OpenStudy (anonymous):

so basically, the two extra forces on lower block will be the friction and normal force from upper block

OpenStudy (anonymous):

with the friction directed in the opp. direction to that acting on upper block

OpenStudy (anonymous):

ok yea I thought the frictional force was probably one. So I add the normal force too even though it acts perpendicular to the motion?

OpenStudy (anonymous):

yes. only because its an inclined surface

OpenStudy (anonymous):

but if you're considering the normal force, you can't write (m1+m2)a=force

OpenStudy (anonymous):

ok so I am kind of confused. Let's pretend that the top block isn't on there. Why wouldn't I also include the normal force of block m2 in my above equation?

OpenStudy (anonymous):

because its perpendicular to the direction of acceleration. but in this case the normal force of block 1 affects friction. it wouldn't however be included in the motion along plane

OpenStudy (anonymous):

the hint is friction always opposes relative motion.